SubjectCover

Question 81 of 108 in Unit 10

Handling Data 7

2 marks · Normal

A bag contains 4 red marbles and 6 blue marbles. A marble is taken at random, replaced, and then another marble is taken at random.

Work out the probability that both marbles are blue.

Work this out on paper, then compare with the answer and worked solution below. With a free account you can submit written answers for marking.

Show answer

0.36 (±0.01)

Also accept: 9/25

Common mistakes: Using 'without replacement' method

Show mark scheme and worked solution

Marking points

M1 · 1 mark · Find probability of blue marble on one pick: 6/10 or 0.6

A1 · 1 mark · Multiply probabilities: 0.6 × 0.6 = 0.36 or 9/25

Worked solution

Probability of picking a blue marble on each draw:

Events are independent (since marble is replaced).

Report a problem with this question
What's wrong?